# HNN extensions of free groups with equal associated subgroups of finite index: polynomial time word problem

### v1

@@ 5. The case of equal associated subgroups of finite index
5.1. $\varphi$-stable subgroups
Let $H\leq F$ and $\varphi\in\operatorname{{Aut}}(H)$. We say that $H^{\prime}\leq H$ is $\varphi$-stable if $\varphi(H^{\prime})\subseteq H^{\prime}$, i.e., if $\varphi|_{H^{\prime}}\in\operatorname{{Aut}}(H^{\prime})$.
Lemma 5.1 (Join).
If $H^{\prime},H^{\prime\prime}$ are $\varphi$-stable, then $\left\langle H^{\prime}\cup H^{\prime\prime}\right\rangle$} is $\varphi$-stable.
@@ 5.1. $\varphi$-stable subgroups
For any $g\in H^{\prime}\cup H^{\prime\prime}$ we have $\varphi(g)\in\varphi(H^{\prime})\cup\varphi(H^{\prime\prime})=H^{\prime}\cup H^{\prime\prime}$. Hence, $\varphi(H^{\prime}\cup H^{\prime\prime})\subseteq H^{\prime}\cup H^{\prime\prime}$. Therefore, $\varphi({\left\langle H^{\prime}\cup H^{\prime\prime}\right\rangle})\subseteq{\left\langle H^{\prime}\cup H^{\prime\prime}\right\rangle}$. ∎
Lemma 5.2 (Meet).
If $H^{\prime},H^{\prime\prime}$ are $\varphi$-stable, then $H^{\prime}\cap H^{\prime\prime}$ is $\varphi$-stable.
@@
For any $g\in H^{\prime}\cap H^{\prime\prime}$ we have $\varphi(g)\in\varphi(H^{\prime})\cap\varphi(H^{\prime\prime})=H^{\prime}\cap H^{\prime\prime}$. Hence, $\varphi(H^{\prime}\cap H^{\prime\prime})\subseteq H^{\prime}\cap H^{\prime\prime}$. ∎
Thus, the set of all $\varphi$-stable subgroups, denoted by $L_{\varphi}$, has a structure of a bounded lattice, with the maximum element $H$ and the minimum element $\{1\}$. Let us consider a set
@@
5.2. $\varphi$-stable normal subgroups of finite index
@@
We suspect that the answer is negative in general. However, how can one find such a subgroup $N$ if it exists? Let us review the properties of a required subgroup $N$. It should satisfy the following four properties:
@@
(6) $\begin{array}[]{l}H_{0}=H,\\ H_{i+1}=\cap_{c}c^{-1}H_{i}c\cap\varphi(H_{i})\cap\varphi^{-1}(H_{i})\ \ \ \mbox{ for }i\geq 0.\end{array}$
@@
Denote $\cap_{i}H_{i}$ by $L$. It is easy to see that it is a normal subgroup of $F$. Also
$x\in L$ $\ \ \Rightarrow\ \ x\in H_{i}\ \forall i$
$\ \ \Rightarrow\ \ \varphi(x)\in\varphi(H_{i})\subseteq H_{i+1}\ \forall i$
@@
Hence, $L\in L_{\varphi}^{\ast}$. By Lemma 5.5, $M_{\varphi}\leq L$. Thus, $L=M_{\varphi}$. ∎
@@
$\varphi$ is normalizable if and only if $\{H_{i}\}$ stabilizes.
@@
$\Leftarrow$” Lemma 5.7.
$\Rightarrow$” If $F:M_{\varphi}]=m<\infty$, then $H_{\lceil\log_{2}(m)\rceil}=M_{\varphi}$ because $H_{i}:H_{i+1}]\geq 2$ whenever $H_{i+1}\neq H_{i}$. ∎
@@ 5.3. The case when $\varphi$ extends to an automorphism of $F$
Recall that a subgroup $N\leq F$ is characteristic if $\varphi(N)\subseteq N$ for every $\varphi\in\operatorname{{Aut}}(F)$. It is easy to see that every characteristic subgroup is normal.
@@ 5.4.1. Syllables
Let $\Gamma^{\ast}=(V,E^{\pm},r,\mu)$ be the subgroup graph for $N$. Fix a set of edges $T\subseteq E^{+}$ defining a spanning tree in $\Gamma$ as described in Section 2.3. The graph $\Gamma^{\ast}$ is the Cayley graph of the finite group $F/N$, which induces a natural group operation on its vertex set $V$. Moreover, since $N\leq H\leq F$, it follows that $H/N\leq F/N$. Hence, for $v\in V$ we write $v\in H$ whenever the coset corresponding to $v$ belongs to $H/N$. In particular, we have $r\in H$.
@@ 5.4.2. Precomputed data
  • For every $v_{1},v_{2}\in V$ we have an SLP $C_{v_{1},v_{2}}=S_{r,v_{1}}(T_{r,v_{2}})\circ T_{v_{1}v_{2},r}$.
@@ 5.4.6. The algorithm
Suppose that $H_{1}=H_{2}$ are finite index subgroups of $F$ and let $\varphi:H_{1}\to H_{2}$ be normalizable. Then the word problem for the HNN extension $F\ast_{\varphi}t$ is decidable in polynomial time.

### v2

@@ 5. The case of equal associated subgroups of finite index
5.1. $\varphi$-invariant subgroups
Let $H\leq F$ and $\varphi\in\operatorname{{Aut}}(H)$. We say that $H^{\prime}\leq H$ is $\varphi$-invariant if $\varphi(H^{\prime})=H^{\prime}$, i.e., if $\varphi|_{H^{\prime}}\in\operatorname{{Aut}}(H^{\prime})$.
Lemma 5.1.
$H^{\prime}\leq H$ is $\varphi$-invariant if and only if $\varphi(H^{\prime})\leq H^{\prime}$ and $\varphi^{-1}(H^{\prime})\leq H^{\prime}$.
@@ 5.1. $\varphi$-invariant subgroups
Applying $\varphi$ to $\varphi^{-1}(H^{\prime})\leq H^{\prime}$, we obtain $H^{\prime}\leq\varphi(H^{\prime})$. Together with the assumption $\varphi(H^{\prime})\leq H^{\prime}$ this implies that $H^{\prime}=\varphi(H^{\prime})$. ∎
Lemma 5.2.
If $\varphi(H^{\prime})\leq H^{\prime}$ and $\varphi(H^{\prime\prime})\leq H^{\prime\prime}$, then
  • (a)
    $\varphi({\left\langle H^{\prime}\cup H^{\prime\prime}\right\rangle})\leq{\left\langle H^{\prime}\cup H^{\prime\prime}\right\rangle}$, and
  • (b)
    $\varphi(H^{\prime}\cap H^{\prime\prime})\leq H^{\prime}\cap H^{\prime\prime}$.
@@
Indeed, for any $g\in F$
[…]
$\ \Rightarrow\ \varphi(g)\in\varphi(H^{\prime})\cup\varphi(H^{\prime\prime})\subseteq H^{\prime}\cup H^{\prime\prime},$
[…]
$\ \Rightarrow\ \varphi(g)\in\varphi(H^{\prime})\cap\varphi(H^{\prime\prime})\leq H^{\prime}\cap H^{\prime\prime},$
which implies (a) and (b). ∎
Lemma 5.3 (Join and meet operations).
If $H^{\prime},H^{\prime\prime}$ are $\varphi$-invariant, then
  • (a)
    $\left\langle H^{\prime}\cup H^{\prime\prime}\right\rangle$} is $\varphi$-invariant, and
  • (b)
    $H^{\prime}\cap H^{\prime\prime}$ is $\varphi$-invariant.
Proof.
Follows from Lemma 5.1 and Lemma 5.2 applied to $\varphi$ and $\varphi^{-1}$. ∎
Thus, the set of all $\varphi$-invariant subgroups, denoted by $L_{\varphi}$, has a structure of a bounded lattice, with the maximum element $H$ and the minimum element $\{1\}$. Let us consider the set
@@
5.2. $\varphi$-invariant normal subgroups of finite index
@@
We suspect that the answer is negative in general. However, how can one find such a subgroup $N$ if it exists? Let us review the properties of a required subgroup $N$. It should satisfy the following three properties:
@@
(6) $\begin{array}[]{l}H_{0}=H,\\ H_{i+1}=\bigcap_{c}c^{-1}H_{i}c\ \cap\ \varphi(H_{i})\ \cap\ \varphi^{-1}(H_{i})\ \ \ \mbox{ for }i\geq 0.\end{array}$
@@
Denote $\cap_{i}H_{i}$ by $L$. It is easy to see that it is a normal subgroup of $F$. Moreover,
$x\in L$ $\ \ \Rightarrow\ \ x\in H_{i}\leq\varphi^{-1}(H_{i-1})\ \forall i\geq 1$
$\ \ \Rightarrow\ \ \varphi(x)\in H_{i-1}\ \forall i\geq 1$
@@
Hence, $\varphi(L)\leq L$ and, similarly, $\varphi^{-1}(L)\leq L$. Therefore, $L$ is $\varphi$-invariant and $L\in L_{\varphi}^{\ast}$. By Lemma 5.6, we have $M_{\varphi}\leq L$. Thus, $L=M_{\varphi}$. ∎
@@
$\varphi$ is normalizable if and only if the sequence $H_{0}\geq H_{1}\geq H_{2}\geq\dots$ stabilizes.
@@
The direction $\Leftarrow$” follows directly from Lemmas 5.5 and 5.8.
To prove the converse, assume that $\varphi$ is normalizable. Then $F:M_{\varphi}]=m<\infty$ and
$H_{\lceil\log_{2}m\rceil}=M_{\varphi},$
since $H_{i}:H_{i+1}]\geq 2$ whenever $H_{i+1}\neq H_{i}$. ∎
@@ 5.3. The case when $\varphi$ extends to an automorphism of $F$
Recall that a subgroup $N\leq F$ is characteristic if $\varphi(N)\leq N$ for every $\varphi\in\operatorname{{Aut}}(F)$. It is easy to see that every characteristic subgroup is normal.
@@ 5.4.1. Syllables
Let $\Gamma^{\ast}=(V,E^{\pm},r,\mu)$ be the subgroup graph for $N$. The graph $\Gamma^{\ast}$ is the Cayley graph of the finite group $F/N$, which induces a natural group operation on its vertex set $V$. Moreover, since $N\leq H\leq F$, it follows that $H/N\leq F/N$. Hence, for $v\in V$ we write $v\in H$ whenever the coset corresponding to $v$ belongs to $H/N$. In particular, we have $r\in H$.
@@ 5.4.2. Precomputed data
  • For every $v_{1},v_{2}\in V$ we have an SLP $C_{v_{1},v_{2}}=S_{r,v_{1}}(T_{r,v_{2}})\circ T_{v_{1}v_{2},r}$, where $v_{1}v_{2}$ in $T_{v_{1}v_{2},r}$ is the product of $v_{1}$ and $v_{2}$ in $F/N$.
@@ 5.4.6. The algorithm
Suppose that $H$ is a subgroup of $F$ of finite index and let $\varphi:H\to H$ be a normalizable isomorphism. Then the word problem for the HNN extension $F\ast_{\varphi}t$ is decidable in polynomial time.